# Quantifying the Effects of Parameters in Widespread SEP Events with EPREM

### v2

@@ 3.4 The Role of Shock Profile
Figures 18 and 19 show proton fluxes from the simulation run with $\Gamma_{s}=10$. The major effect of softening the shock-front gradient is to reduce fluxes below their baseline values at all observers, with the amount of reduction increasing with energy. This broad reduction resulted in the observers located $180^{\circ}$ from the shock origin recording only negligible flux in the final day of simulation time. The softer shock front also causes fluxes observed at locations $0^{\circ}$$90^{\circ}$ from the shock origin to return to their pre-shock values more quickly. West of the shock origin, the flux of protons with $\textMeV\leq 10$$\text${MeV}} is more distinctly separated into a band at 500 keV–1 MeV, and one at 5 MeV–10 MeV.
@@
4 Discussion
[…]
@@ 4 Discussion
Changing the dependence of mean free path on magnetic-field strength and rigidity had the least noticeable effect on fluxes observed at $180^{\circ}$ from the shock origin, especially for protons with 50 and 100 MeV at 1.0 au. However, the values of those parameters strongly affected timing of peak fluxes and the evolution of flux as a function of energy at all other observers. This means that simulation runs designed to reproduce proton fluxes observed by spacecraft may need to tune the value of each parameter in order to improve agreement. Quasi-linear theory developed by Jokipii (1966, 1968) and extended by Erdos et al. (1999) with the help of Ulysses observations suggests $\lambda_{\parallel}\left(B\right)\propto|\vec{B}|^{-\beta/2}$ with $2/3\leq\beta\leq 2$. The collection of simulation runs presented here include two values of $\beta$ within that range (namely, 1 and 2) and one value of $\beta$ above the proposed upper bound (namely, 4). The latter value had a drastic effect on the simulated fluxes recorded by point observers, leading to possibly non-physical results, but those results are nonetheless valuable if they can explain unusual in situ observations by spacecraft. Fluxes in a simulation run with $\beta=2/3$ (not shown) continue two trends seen in the mfp-power-hi, baseline, and mfp-power-lo runs: further decreasing $\beta$ leads to 1) further reduced fluxes at energies $\textMeV>1$$\text${MeV}} and 2) a stronger pre-shock ramp-up in fluxes at energies $\textMeV\leq 1$$\text${MeV}}. Dröge (2000) derived estimates of the mean free path dependence on rigidity by computing fits to the ISEE-3/ICE spacecraft. They found a consistent power-law index of roughly 0.3 for ions despite large variation in mean-free-path amplitude, which ranged from 0.02 au to more that 1.0 au. EPREM is not currently capable of running with $\chi=0$ (i.e., a mean free path that is constant with respect to rigidity) because the lower bound is set at $10^{-33}$. However, fluxes from a run with $\chi=10^{-33}$ (not shown) continued the main trends between $\chi=2/3$ in the mfp-rigidity run to $\chi=1/3$ in the baseline run: higher flux of protons with energy $\textMeV\leq 1$$\text${MeV}}, lower flux of protons with energy $\textMeV\geq 50$$\text${MeV}}, and flux increases that occur essentially contemporaneously across energies east of the shock origin.
@@ 5 Concluding Remarks
The authors acknowledge extensive contributions by Ronald Caplan to the EPREM source code, especially to the algorithms used to solve the focused transport equation. This work benefited from previous conversations with Christina Cohen, Gabriel Muro, and Erika Palmerio about spacecraft observations of specific widespread SEP events. {contribution}MY developed the analysis code, maintains the EPREM-related code repositories, performed the simulation runs, created the figures, and wrote the original manuscript. BP provided the funding and edited the manuscript. Both authors were responsible for developing the initial research concept.
@@ Appendix B Lagrangian Grid
[…]
@@ Appendix C Ideal Shock
[…]

### v3

@@ III.4 The Role of Shock Profile
Figures 18 and 19 show proton fluxes from the simulation run with $\Gamma_{s}=10$. The major effect of softening the shock-front gradient is to reduce fluxes below their baseline values at all observers, with the amount of reduction increasing with energy. This broad reduction resulted in the observers located $180^{\circ}$ from the shock origin recording only negligible flux in the final day of simulation time. The softer shock front also causes fluxes observed at locations $0^{\circ}$$90^{\circ}$ from the shock origin to return to their pre-shock values more quickly. West of the shock origin, the flux of protons with $\leq 10{\text{ MeV}}$ is more distinctly separated into a band at 500 keV–1 MeV, and one at 5 MeV–10 MeV.
@@
IV Discussion
Despite EPREM’s simple built-in shock model, the simulation runs presented in §III can tell us a good deal about the nature of widespread SEP events. First of all, it is not difficult to produce flux enhancements at energies up to 100 MeV over a span of $\pm 90^{\circ}$ from the shock origin, though the ability to observe these flux enhancements depends not only on an observer’s location but also on the sensitivity of a given instrument. Assuming a typical noise floor of $10^{-3}{\text{ counts}}{\text{ cm}}^{-2}{\text{ s}}^{-1}{\text{ sr}}^{-1}{\text{ (MeV/nuc)}}^{-1}$ (e.g., Kollhoff et al., 2021; Dresing et al., 2023), most simulation runs shown here generated observable flux of 100-MeV protons at the location of at least one observer. Of those that caused one or more observers to not observe 100-MeV protons above this fiducial threshold, increasing the reference parallel mean free path, $\lambda_{\parallel 0}$, played the most significant role (cf. Figure 7), followed by weakening the shock gradient (cf. Figure 18). Setting $\lambda_{\parallel}\left(r\right)\propto r^{-2}$ everywhere (cf. Figure 15) preferentially reduced 100-MeV fluxes west of the shock origin while setting $\lambda_{\parallel}\left(|\vec{B}|\right)\propto|\vec{B}|^{-1/2}$ preferentially reduced 100-MeV fluxes at and east of the shock origin.
We can test the general validity of our baseline simulation run by comparing the fluxes shown in Figure 3 to those shown in the panel of Figure 1 in Dresing et al. (2023). During the 17 April 2021 widespread SEP event, BepiColombo was best connected to the flare site, at a radial distance of 0.63 au, and therefore best corresponds to our point observer 00. It observed a sharp increase in the flux of 25.1 MeV protons roughly one hour after the associated flaring event, peaking at approximately $4-5\times 10^{-1}{\text{ counts}}{\text{ cm}}^{-2}{\text{ s}}^{-1}{\text{ sr}}^{-1}{\text{ (MeV/nuc)}}^{-1}$, then exponentially decaying to the background level over the following 30 hours. The baseline flux of 10-MeV protons recorded by observer 00 qualitatively mirrors this behavior; in fact, assuming the peak flux of 25.1 MeV protons should lie between those of 10 MeV and 50 MeV protons, quantitatively agrees reasonably well. Point observer 07 recorded a slightly lower peak flux and faster decay, similar to that observed by Parker Solar Probe, while point observers 02 and 03 recorded flatter profiles with still lower peaks, similar to observed by Solar Orbiter. There is even some consistency between point observer 11 and STEREO-A, which observed weak fluxes of protons with 20.8 to 23.8 MeV, beginning later than the peak intensity observed by BepiColombo and Parker Solar Probe.
@@ IV Discussion
The mean-free-path inverse dependence on $|\vec{B}|$ also affects diffusion. On one hand, streams from run mfp-power-hi show larger fluxes of 100-MeV protons near $180^{\circ}$ with broader radial spread. This is consistent with the mean free path growing even more quickly with radial distance, which allows these protons to diffuse outward more quickly, and can thus account for the earlier reduction in 100-MeV protons recorded by observer 04 (at $180^{\circ}$ and 0.5 au). On the other hand, observer 04 recorded a slight increase in lower-energy protons during run mfp-power-lo, for which the mean free path had the least dependence on $|\vec{B}|$ of all simulation runs shown here. This reduced dependence on $|\vec{B}|$ leads to reduced perpendicular diffusion at all energies, compared to baseline values. However, it also reduces the maximum energy that protons acquire before leaving the acceleration region (Schwadron et al., 2015). We can therefore attribute the slight increase in 1-MeV protons recorded by observer 04 to diffusion from regions near observers 05 and 13 (i.e., far east of the shock origin).
Changing the dependence of mean free path on magnetic-field strength and rigidity had the least noticeable effect on fluxes observed at $180^{\circ}$ from the shock origin, especially for protons with 50 and 100 MeV at 1.0 au. However, the values of those parameters strongly affected timing of peak fluxes and the evolution of flux as a function of energy at all other observers. This means that simulation runs designed to reproduce proton fluxes observed by spacecraft may need to tune the value of each parameter in order to improve agreement. Quasi-linear theory developed by Jokipii (1966, 1968) and extended by Erdos et al. (1999) with the help of Ulysses observations suggests $\lambda_{\parallel}\left(B\right)\propto|\vec{B}|^{-\beta/2}$ with $2/3\leq\beta\leq 2$. The collection of simulation runs presented here include two values of $\beta$ within that range (namely, 1 and 2) and one value of $\beta$ above the proposed upper bound (namely, 4). The latter value had a drastic effect on the simulated fluxes recorded by point observers, leading to possibly non-physical results, but those results are nonetheless valuable if they can explain unusual in situ observations by spacecraft. Fluxes in a simulation run with $\beta=2/3$ (not shown) continue two trends seen in the mfp-power-hi, baseline, and mfp-power-lo runs: further decreasing $\beta$ leads to 1) further reduced fluxes at energies $1{\text{ MeV}}$ and 2) a stronger pre-shock ramp-up in fluxes at energies $\leq 1{\text{ MeV}}$. Dröge (2000) derived estimates of the mean free path dependence on rigidity by computing fits to the ISEE-3/ICE spacecraft. They found a consistent power-law index of roughly 0.3 for ions despite large variation in mean-free-path amplitude, which ranged from 0.02 au to more that 1.0 au. EPREM is not currently capable of running with $\chi=0$ (i.e., a mean free path that is constant with respect to rigidity) because the lower bound is set at $10^{-33}$. However, fluxes from a run with $\chi=10^{-33}$ (not shown) continued the main trends between $\chi=2/3$ in the mfp-rigidity run to $\chi=1/3$ in the baseline run: higher flux of protons with energy $\leq 1{\text{ MeV}}$, lower flux of protons with energy $\geq 50{\text{ MeV}}$, and flux increases that occur essentially contemporaneously across energies east of the shock origin.
@@ V Concluding Remarks
The authors acknowledge extensive contributions by Ronald Caplan to the EPREM source code, especially to the algorithms used to solve the focused transport equation. This work benefited from previous conversations with Christina Cohen, Gabriel Muro, and Erika Palmerio about spacecraft observations of specific widespread SEP events.
{contribution}
MY developed the analysis code, maintains the EPREM-related code repositories, performed the simulation runs, created the figures, and wrote the original manuscript. BP provided the funding and edited the manuscript. Both authors were responsible for developing the initial research concept.
@@ Appendix B Lagrangian Grid
EPREM manages a set of surfaces on which it places nodes for calculation. There are six computational surfaces that are logically arranged as the faces of a cube. However, the edges of these faces define logical connections only — not rigid boundaries (e.g., as on a Euclidian cube). The outward movement of computational nodes effectively deforms the cube of logical faces into a cubed sphere (e.g., Figure 1 of Sadourny, 1972). Each node advances according to $\delta\vec{r}=\vec{V}\delta t$, where $\vec{r}$ is the 3D node displacement, $\vec{V}$ is the 3D flow velocity, and $\delta t$ is the time step duration. It then calculates $f_{s}$ according to the FTE. Each set of linked nodes (typically called a “stream”) defines a velocity path line that tracks the trajectory of the distribution of particles. In purely steady-state conditions, these velocity path lines represent formal streamlines. If, and only if, the frozen-in condition of MHD holds, they also represent magnetic field lines. EPREM creates the initial grid (known as “seeding” the nodes) by generating a new node on every stream, pushing the node to $r=V_{r}\delta t$, and rotating the inner boundary by $\Delta\phi=\Omega_{\odot}\delta t$, where $V_{r}$ is the radial component of the solar wind and $\Omega_{\odot}$ is the solar rotation rate. It repeats this process until it has seeded all nodes. The default configuration of EPREM streams is therefore a Parker spiral with the outermost node at $R_{\text{max}}=R_{\text{min}}+NV_{r}\delta t$, where $R_{\text{min}}$ is the location of the inner boundary (typically the solar surface) and $N$ is the number of nodes. By nature of the Lagrangian grid, EPREM does not have a fixed outer boundary.
@@ Appendix C Ideal Shock
The EPREM ideal shock model modifies the default magnetic-field components, velocity-field components, and density to simulate a simple radially propagating shock. The density jump of an EPREM ideal shock follows the standard Rankine-Hugoniot relation $n_{1}=q_{s}n_{2}$, where subscripts 1 and 2 denote downstream and upstream quantities, respectively, and $q_{s}$ is the compression ratio. The form of the background (i.e., unshocked) density is $n(r)=n_{0}\left(\text{1 au}/r\right)^{2}$, where $n_{0}$ is the density at 1 au. The user may set $q_{s}$ and $n_{0}$ at runtime.